# Hamiltonian stochastic Lie systems and applications

### v2

@@ Acknowledgements
Fruitful discussions with J.P. Ortega about setting a correction for his stochastic Lie theorem are acknowledged. X. Rivas acknowledges partial financial support from the Spanish Ministry of Science and Innovation grants PID2021-125515NB-C21 and RED2022-134301-T of AEI, and the Ministry of Research and Universities of the Catalan Government project 2021 SGR 00603. J. de Lucas acknowledges a MINIATURA-5 grant by the Polish Narodowe Centrum Nauki under project Nr 2021/05/X/ST1/01797. The research of Marcin Zając was funded by the Polish National Science Centre under the contract number UMO-2021/41/N/ST1/02908. For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. The authors also acknowledge the contribution of RED2022-134301-T funded by MCIN/AEI/10.13039/501100011033 (Spain).

### v3

@@ Acknowledgements
We acknowledge fruitful discussions with J. P. Ortega on establishing a correction for his stochastic Lie theorem. We express our gratitude to two anonymous referees for their valuable feedback, which motivated us to enhance our manuscript considerably by refining and expanding examples, adding new insights into our theoretical framework, and offering suggestions to make it more understandable to a broader readership. X. Rivas acknowledges partial financial support from the Spanish Ministry of Science and Innovation grants PID2021-125515NB-C21 and RED2022-134301-T of AEI, and the Ministry of Research and Universities of the Catalan Government project 2021 SGR 00603. The research of Marcin Zając was funded by the Polish National Science Centre under a PRELUDIUM project with contract number UMO-2021/41/N/ST1/02908. For the purpose of Open Access, the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. The authors also acknowledge the contribution of RED2022-134301-T funded by MCIN/AEI/10.13039/501100011033 (Spain).